the equation of the tangent to the circle x2 +y2 +4x -4y+4=0 which make equal intercept on the positive coordinate axes is:

Grade:12

1 Answers

Vaishal Shah
12 years ago
if tangent is y=mx+c. Equal intercept means m=-1 x+y=c perpendicular from center of the circle on tangent is radius. C=2 root 2 x+y=2 root 2